Solve differential equation

(D^3-6D’D+11D’^2D-6D’)z=cos(2x+y)+e^(2x+y)-y


1
Expert's answer
2021-12-06T16:49:25-0500

"(D^3-6D\u2019D+11D\u2019^2D-6D\u2019)z=cos(2x+y)+e^{2x+y}-y"


auxillary equation:

"m^3-6m+11m-6=m^3+5m-6=0"

"(m-1)(m^2+m+6)=0"


complementary function:

"C.F.=f_1(y+x)+\sum c_re^{a_rx+b_ry}"

where

"a_r^2+b_r+6=0"

then:

"C.F.=f_1(y+x)+\sum c_re^{a_rx-(6+a_r^2)y}=f_1(y+x)+e^{-6y}\sum c_re^{a_r(x-a_ry)}"


particular Integral:

"P.I._1=\frac{e^{2x+y}}{D^3-6D’D+11D’^2D-6D’}" (replace D by 2 and D' by 1)


"=\frac{e^{2x+y}}{8-12+22-6}=\frac{e^{2x+y}}{12}"


"P.I._2=\frac{cos(2x+y)}{D^3-6D’D+11D’^2D-6D’}" (replace D2=-a2=-4,DD'=-ab=-2,D'2=-b2=-1)


"=\frac{cos(2x+y)}{-4D+12-11D-6D'}=\frac{cos(2x+y)}{-15D+12-6D'}=-\frac{cos(2x+y)/15}{D+0.4D'-0.8}"


"\frac{1}{D-mD'}F(x,y)=\int F(x,c-mx)dx"


where c is replaced by y+mx after integration

then:


"P.I._2=-\frac{cos(2x+y)/15}{D+0.4D'-0.8}=-\frac{1}{15}\intop cos(2x+c+0.4x)dx"


"=-\frac{sin(2.4x+c)}{15\cdot2.4}=-\frac{sin(2x+y)}{36}"


"P.I._3=\frac{-y}{D^3-6D’D+11D’^2D-6D’}=-\frac{1}{D^3}(1-\frac{6D'}{D^2}+\frac{11D'^2}{D^2}-\frac{D'}{d^3})^{-1}y="


"=-\frac{1}{D^3}(1+\frac{6D'}{D^2}-...)y=-\frac{1}{D^3}(y+\frac{6}{D^2})=-\frac{1}{D^3}(y+3x^2)=-(yx^3/3+x^5/20)"


"z=C.F.+P.I.="


"=f_1(y+x)+e^{-6y}\sum c_re^{a_r(x-a_ry)}+\frac{e^{2x+y}}{12}-\frac{sin(2x+y)}{36}-(yx^3/3+x^5/20)"